OpenStudy (anonymous):

Shift the indice in the summation:

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}a _{k}x ^{3k}\]

OpenStudy (anonymous):

I need x^3k to be x^k but I am drawing a blank on how to shift this correctly!

OpenStudy (whpalmer4):

Does it help to rewrite \[x^{3k}= x^k*x^k*x^k\]? also, are you sure you have written that properly? you used \(n\) on the summation sign, but \(k\) elsewhere...

OpenStudy (anonymous):

Oops, I meant k=0....

OpenStudy (anonymous):

hmm not sure if that would help... Basically I have simplified my question from what originally was asking for me to solve a diff eqs using power series and I came across a situation where I needed to shift indices to collect terms together.

OpenStudy (anonymous):

I could write out the whole thing and maybe you will see better than I...

OpenStudy (whpalmer4):

I don't see how you can shift \[a_0 x^0 + a_1 x^3+ a_2 x^6 + ... a_kx^{3k} \] into something of the form \[x^0+x^1+x^2+x^3 + ... + x^k\]

OpenStudy (anonymous):

I must have made a mistake somewhere then... here is the whole darn thing:..

OpenStudy (anonymous):

\[\sum_{n=2}^{\infty}(n-1)na _{n} x ^{n-2}+\sum_{n=0}^{\infty}(-1)^{n}x ^{2n}\sum_{n=0}^{\infty}a _{n} x ^{n}\]

OpenStudy (anonymous):

\[\sum_{k=0}^{\infty}(k+1)(K+2)a _{k+2} x ^{K}+\sum_{k=0}^{\infty}(-1)^{k}x ^{3k}a ^{k}\]

OpenStudy (anonymous):

so I combined those last two summations for a x^3k but need a x^k

OpenStudy (anonymous):

****also forgot to put in the divide by (2k)! on that last summation...****

OpenStudy (anonymous):

that might be where the trick is...

OpenStudy (anonymous):

\[\[\sum_{k=0}^{\infty}\frac{ (-1)^{k}x ^{3k}a _{k} }{ (2k)!}\\]\]

OpenStudy (amistre64):

to shift stuff adding to the index subtracts from the rule multiplying to the index divides from the rule in effect, you are trying to keep the terms of the summation the same regardless of the indexing

OpenStudy (anonymous):

terminology was wrong. Thanks for the correction. How would I get the term x^3k to be just x^k...

OpenStudy (amistre64):

\[\sum_{k=0}^{\infty}a _{k}x ^{3k}=a_0x^0+a_1x^3+a^2x^6+...\] \[\sum_{k/3=0}^{\infty}a _{(k/3)}x ^{3(k/3)}=a_0x^0+a_1x^3+a^2x^6+...\] \[\large\sum_{k=0*3}^{\infty}a _{(k/3)}x ^{k}\]

OpenStudy (amistre64):

also, x^(3k) = (x^3)^k

OpenStudy (anonymous):

the a^2 should be a _{2}, am I right?

OpenStudy (amistre64):

lol, yes ...

OpenStudy (amistre64):

but from the top posts, can you explain how you resolve a product of sums?

OpenStudy (anonymous):

stupid question I know but I am a little baffled by the k/2=0 indice because I thought you could only have whole numbers for a...

OpenStudy (anonymous):

what doe you mean resolve ? Simplify? I thought I had done it right..

OpenStudy (amistre64):

lets observe a shift by adding \[\Large\sum_{k=0}^{k=8}a_kx^{k-1}\]add 1 to each k \[\Large\sum_{(k+1)=0}^{(k+1)=8}a_{(k+1)}x^{(k+1)-1}\]and simplify \[\Large\sum_{k=-1}^{k=7}a_{k+1}x^{k}\]

OpenStudy (anonymous):

yes, this makes sense to me

OpenStudy (amistre64):

lets try this by modifying k by a factor \[\Large\sum_{k=0}^{k=8}a_kx^{3k}\] divide each k by 3 \[\Large\sum_{k/3=0}^{k/3=8}a_{k/3}x^{3(k/3)}\]and simplify \[\Large\sum_{k=0}^{k=24}a_{k/3}x^{k}\] since you only want the integer values of k/3, you might have to consider removing some fractioanl aks

OpenStudy (amistre64):

a0 a1/3 x a2/3 x^2 a3/3 x^3 = a1 x^3

OpenStudy (anonymous):

voilà !!

OpenStudy (anonymous):

wait..... that last part got me...with the a's again..

OpenStudy (anonymous):

a _{1/3} <--how can that be?

OpenStudy (anonymous):

I mean that is perfectly fine in summations? no rules against that?

OpenStudy (amistre64):

its still a discrete count, but its not really a part of the original setup so it might have to be excluded is all

OpenStudy (amistre64):

the factor produces a unit length of 1.3 instead of 1 is all

OpenStudy (amistre64):

*1/3

OpenStudy (amistre64):